# Trust Me, I’m Wrong: High-Certainty Hallucinations in LLMs

### v1

@@ Appendix G CHOKE Uniqueness – Additional Results
Tables 6 and 7 display these results. Notably, all values are higher than those reported in the main paper, as the examples are now sampled from a subset of shared hallucination examples between the Child and Alice-Bob settings. However, even under the shared examples, the Jaccard similarity for certain cases remains significantly higher than the similarity observed under the permutation test. This further supports the conclusion that certain hallucinations are not mere noise but a distinct property of the models.
We compare the similarity under high certainty to that under the lowest certainty. Specifically, we examine the same subset of examples as the high-certainty subgroup but focus on those with the lowest probabilities. The results are shown in Table 8.
@@
Model Dataset Random Certain
@@
Table 6: Jaccard Similarity of CHOKE hallucinations across different prompts under shared hallucinations. The Certain column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets, using Probability as the certainty threshold. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant ($p<0.0001$, permutation test (the Rand column).
Model Dataset Random Certain
@@
Table 7: Jaccard Similarity of CHOKE hallucinations across different prompts under shared hallucinations. The Certain column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets, using Semantic Entropy as the certainty threshold. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant ($p<0.0001$, permutation test (the Rand column).
Model Dataset uncertain Certain
@@
Table 8: Jaccard Similarity of CHOKE hallucinations across different prompts. The Certain column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets and Low certain shows the results of the lowest certainty subset, using Probability as the certainty threshold. Results indicate high similarity, suggesting consistency across settings.
@@ Appendix H CHOKE Persists in Larger Models – Additional Results
In Section 4.2, we demonstrated that Gemma-2-27B achieves similar or slightly higher CHOKE detection results on the TriviaQA dataset than Gemma-2-9B, thus showing the existence of this phenomenon on larger models. To further illustrate this phenomenon, Figure 10 presents comparable results on the Natural Questions dataset. These results show a clear correlation with those presented in the main paper.
(a) Gemma, Child
(b) Gemma, Alice-Bob
Figure 10: Detection of CHOKE Hallucinations: Comparing Gemma-27B to Gemma-9B CHOKE hallucinations on Natural Questions. The results indicate that CHOKE hallucinations is similar and slightly higher for the larger model, Gemma-27B.

### v2

@@ Appendix G CHOKE Uniqueness – Additional Results
We start by showing that using TriviaQA we get similar results to the ones in the main paper using Natural Questions. See Table 8. Those results show the consistency of the CHOKE uniqueness across datasets.
Next, we extend our evaluation to shared hallucination examples using an analysis on two settings (4,5). Tables 9 and 10 display these results on probability and Semantic entropy. Notably, the values are higher than those reported in the main paper, as the examples are now sampled from a subset of shared hallucination examples between two settings. However, even under the shared examples, the Jaccard similarity for certain cases remains significantly higher than the similarity observed under the permutation test. This further supports the conclusion that certain hallucinations are not mere noise but a distinct property of the models.
Lastly, we compare the similarity under high certainty to that under the lowest certainty. Specifically, we examine the same subset of examples as the high-certainty subgroup but focus on those with the lowest probabilities. The results are shown in Table 11.
@@
Semantic Entropy Probability
Model Random CHOKE Random CHOKE
Llama $3.3{\pm 1.1}$ $\mathbf{18.4}{\pm 5.6}$ $3.2{\pm 0.9}$ $\mathbf{30.3}{\pm 14.3}$
Mistral $3.4{\pm 1.0}$ $\mathbf{15.4}{\pm 4.1}$ $5.5{\pm 1.3}$ $\mathbf{36.6}{\pm 12.3}$
Gemma $3.7{\pm 0.9}$ $\mathbf{17.0}{\pm 8.3}$ $2.9{\pm 0.9}$ $\mathbf{24.9}{\pm 12.9}$
Llama-Inst $4.3{\pm 1.5}$ $\mathbf{19.0}{\pm 7.2}$ $8.2{\pm 2.0}$ $\mathbf{27.0}{\pm 11.5}$
Mistral-Inst $7.4{\pm 0.9}$ $\mathbf{21.6}{\pm 7.0}$ $7.4{\pm 1.8}$ $\mathbf{33.5}{\pm 13.6}$
Gemma-Inst $6.9{\pm 2.0}$ $\mathbf{25.6}{\pm 12.3}$ $7.7{\pm 1.4}$ $\mathbf{32.4}{\pm 17.3}$
Table 8: Jaccard Similarity of CHOKE across different prompts. The CHOKE columns shows the overall similarity of CHOKE samples between prompts in the TriviaQA dataset, using Semantic entropy, and Probability as the certainty thresholds. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant (permutation test, Random column, $p<0.008$).
Model Dataset Random CHOKE
@@
Table 9: Jaccard Similarity of CHOKE hallucinations across different prompts under shared hallucinations. The CHOKE column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets, using Probability as the certainty threshold. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant ($p<0.0001$, permutation test (the Rand column).
Model Dataset Random CHOKE
@@
Table 10: Jaccard Similarity of CHOKE hallucinations across different prompts under shared hallucinations. The CHOKE column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets, using Semantic Entropy as the certainty threshold. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant ($p<0.0001$, permutation test (the Rand column).
Model Dataset uncertain CHOKE
@@
Table 11: Jaccard Similarity of CHOKE hallucinations across different prompts. The CHOKE column shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets and Low certain shows the results of the lowest certainty subset, using Probability as the certainty threshold. Results indicate high similarity, suggesting consistency across settings.
@@ Appendix H CHOKE Persists in Larger Models – Additional Results
In Section 4.2, we demonstrated that Gemma-2-27B achieves similar or slightly higher CHOKE detection results on the Natural Questions dataset than Gemma-2-9B, thus showing the existence of this phenomenon on larger models. To further illustrate this phenomenon, Figure 12 presents comparable results on the TriviaQA dataset. These results show a clear correlation with those presented in the main paper.
Figure 12: Detection of CHOKE: Comparing Gemma-27B to Gemma-9B on TriviaQA. We report both on hallucinations (H) and on non-hallucinations (NH) data. The results indicate that certainty levels are comparable and slightly higher for the larger model, Gemma-27B.
Appendix I CHOKE Score- Additional Results
Similar to the results presented in Section 5, we provide additional evaluations on TriviaQA. See Figure 13 for the full results. As with the Natural Questions setting, we observe that the CHK-F and CHK scores remain lower than the accuracy scores for certainty-based methods, reinforcing the conclusions drawn in the main experiments. This further supports the utility of the CHOKE score as a complementary evaluation measure for mitigation methods, capturing nuances that traditional metrics may miss.
To better emphasize the importance of CHOKE score, we also compare it to F1 and AUROC. [^6] We ran AUROC only on non-prompt methods, as this method requires numerical scores. In Figures 14 and 15, we show for Natural Questions and TriviaQA datasets that those two scores also do not capture what CHK and CHK-F capture, further emphasizing the importance of the new scores.
Figure 13: Our mitigation outperforms on CHOKE-Score and CHOKE-Score reveals limits of standard methods. Averaged over six models and all prompts on TriviaQA, the left figure shows our probe method achieves the highest CHOKE-Score scores. The right figure compares CHOKE-Score (red) to other metrics (blue shades), showing certainty methods perform well generally but poorly on CHOKE-Score, exposing gaps in handling CHOKE hallucinations. Probe methods maintain more consistent performance, demonstrating stronger robustness.
Figure 14: CHOKE-Score Exposes Limitations of Standard Hallucinations Mitigation Methods. Performance of mitigation methods, averaged across six models on Natural Questions. We report AUROC (AUROC), F1 (F1), and the proposed CHOKE-Scores: strict (CHK) and flexible (CHK-F). While certainty and prompt based methods perform well on standard metrics, their CHK scores are substantially lower, revealing a gap in handling CHOKE hallucinations. Probe-based methods, in contrast, maintain consistent performance across all metrics, indicating stronger robustness to CHOKE examples.
Figure 15: CHOKE-Score Exposes Limitations of Standard Hallucinations Mitigation Methods. Performance of mitigation methods, averaged across six models on TriviaQA. We report AUROC (AUROC), F1 (F1), and the proposed CHOKE-Scores: strict (CHK) and flexible (CHK-F). While certainty and prompt based methods perform well on standard metrics, their CHK scores are substantially lower, revealing a gap in handling CHOKE hallucinations. Probe-based methods, in contrast, maintain consistent performance across all metrics, indicating stronger robustness to CHOKE examples.